# Regional and partial observability and control of waves

### v1

@@ 3.8. A $1$-d example
(3.10) $\left\{\begin{array}[]{ll}\partial_{t}^{2}u-\partial_{x}^{2}u=0,&\text{in }\,\,(0,T)\times(-10,10),\\ u(t,-10)=u(t,10)=0,&\text{on }\,\,(0,T),\\ (u(0,\cdot),\partial_{t}u(0,\cdot))=(u_{0},u_{1}),&\end{array}\right.$
@@
(3.11) $\left\{\begin{array}[]{ll}\partial_{t}w_{+}+\partial_{x}w_{+}=0,&\text{in }\,\,(0,T)\times(-10,10),\\ \partial_{t}w_{-}-\partial_{x}w_{-}=0,&\text{in }\,\,(0,T)\times(-10,10),\\ (w_{+}+w_{-})(t,-10)=(w_{+}+w_{-})(t,10)=0,&\text{on }\,\,(0,T),\\ (w_{+}(0,\cdot),w_{-}(0,\cdot))=(u_{1}-\partial_{x}u_{0},u_{1}+\partial_{x}u_{0}),&\end{array}\right.$
In this case, the bicharacteristic rays are particularly simple: they are the curves $t\mapsto x_{0}+t$ and $t\mapsto x_{0}-t$ for $x_{0}\in(-10,10)$ while these curves stay in the domain, bouncing back when meeting the boundary. Accordingly, in $1$-d, we can identify the characteristic manifold $\text{Char}(P_{A})$ with ×${\mathbb{R}}\times\Omega\times\{-1,1\}$, depending if $\tau=|\xi|$, corresponding to $\epsilon=1$, or $\tau=-|\xi|$ corresponding to $\epsilon=-1$.
@@
$\mathscr{R}(\omega_{T})^{+}$ =$=\{(t,x,+1)\text{ s.t. }t\in[0,T]\text{ and }-5+t<x<2+t\}$
$\mathscr{R}(\omega_{T})^{-}$ =$=\{(t,x,-1)\text{ s.t. }t\in[0,T]\text{ and }-2-t<x<5-t\},$
@@ 4. Further observability results
The aim of this section is to refine Theorem 1.1 by analyzing which microlocal components of the initial data of general wave solutions can be effectively observed from measurements taken on $\omega_{T}$.
@@ 4.1. Statement of the results
(4.7) $\|\varphi(x)(u_{0},u_{1})\|_{H^{1}_{0}(\Omega)\times L^{2}(\Omega)}\leqslant C\|\partial_{t}u\|_{L^{2}(\omega_{T})}+C\|(1-\varphi(x))(u_{0},u_{1})\|_{L^{2}(\Omega)\times H^{-1}(\Omega)}.$
Similarly, for any initial data $(u_{0},u_{1})\in L^{2}(\Omega)\times H^{-1}(\Omega)$
(4.8) $\|\varphi(x)(u_{0},u_{1})\|_{L^{2}(\Omega)\times H^{-1}(\Omega)}\leqslant C\|u\|_{L^{2}(\omega_{T})}+C\|(1-\varphi(x))(u_{0},u_{1})\|_{H^{-1}(\Omega)\times(H^{2}\cap H^{1}_{0})^{\prime}(\Omega)}.$
@@
  • The constants appearing in Theorem 4.1 depend on both the observation time $T$ and the underlying metric $(a^{ij}(x))$. While the dependence on $T$ is straightforward, the dependence on the metric is more subtle and difficult to quantify explicitly, as it involves a large number of derivatives of the coefficients $a^{ij}$. This dependence can, in principle, be traced through a careful analysis of the proof of Theorem 3.5 in [15]. However, it is not readily expressible in closed form, since the operator norm of a pseudo-differential operator typically involves multiple derivatives of its symbol, making the exact dependence implicit and technically intricate.
@@
To illustrate the difficulty, consider the inequality (4.1). Let $\psi=\psi(t,x,D_{t},D_{x})\in\mathcal{A}^{0}$ be a pseudodifferential operator with $\operatorname{supp}\psi\cap T^{*}b\mathcal{L}\subset\mathscr{R}(\omega_{T})$, and define the set
@@
By (4.2), any $u\in\mathcal{N}\psi$ satisfies $u=\psi u\in H^{1}_{\text{loc}}(\mathcal{L})$, implying that $\mathcal{N}_{\psi}$ is compact and hence finite-dimensional.
Now consider $v=\partial_{t}u$ for $u\in\mathcal{N}\psi$. Clearly, $v$ also solves the wave equation and satisfies $\partial_{t}v=0$ in $\omega_{T}$. However, there is no guarantee that v belongs to $\mathcal{N}\psi$, as we do not have $(I-\psi)v=0$. In fact, since $\psi v=\psi(\partial_{t}u)$ and $v=\partial_{t}u$, their difference involves the commutator $[\psi,\partial_{t}]u$, which does not vanish in general. Therefore, the naive use of the operator $\partial_{t}$ does not yield an operator acting invariantly on $\mathcal{N}_{\psi}$.
We have not been able to further analyze the structure of the sets $\mathcal{N}\psi$. Whether or not $\mathcal{N}\psi$ is non-trivial remains an open problem.
@@
[…]
@@
Then all the items (1a), (2a), (1b) and (2b) holds for solutions $u$ of the wave equation (1.1) corresponding to initial data $(u_{0},u_{1})$ with either $u_{0}=0$ or $u_{1}=0$.
@@
Figure 4. When $\Omega={\mathbb{R}}^{d}$, $\omega$ is the unit ball and $\Delta_{A}$ is the flat Laplacian, the set $\mathscr{O}(T)$, when projected in the physical space is supported in the ball of size $T$. In fact, in this case, as the bicharacteristics in this setting correspond to straight lines,
$\mathscr{O}(T)=\{(x,\xi)\in T_{b}^{*}\Omega\backslash 0\text{ s.t. }\exists t\in[0,T]\text{ satisfying }x+\xi t\in\omega\}.$
@@ 4.2. Proofs
By contradiction and following the proof of the estimate (4.11) of Theorem 4.7, we get a sequence $u_{1,k}$ such that $((I-\psi(x,D_{x}))u_{1,k})_{k\in{\mathbb{N}}}$ goes to 00 in $H^{-1}(\Omega)$, $\psi(x,D_{x})u_{1,k}$ is of unit norm in $H^{-1}(\Omega)$, and such that the corresponding solutions $u_{k}$ of (1.1) satisfies that $(\partial_{t}u_{k})_{k\in{\mathbb{N}}}$ goes to 00 in $L^{2}(\omega_{T})$. Consequently, up to a subsequence, we get $u_{1}$ such that $(u_{1,k})$ converges weakly to $u_{1}$ in $H^{-1}(\Omega)$, and $(I-\psi(x,D_{x}))u_{1}=0$, and such that the corresponding solution of (1.1) with initial data $(0,u_{1})$ satisfies $\partial_{t}u=0$ in $(0,T)\times\omega$. By the above uniqueness result, we thus get that $u_{1}=0$ in the set $O^{T}=\{x\in\overline{\Omega},\,d(x,\omega)<T\}$. Finally, since $\psi=\psi(x,D_{x})\in\mathcal{B}^{0}$ satisfies $\operatorname{Supp}(\psi)\cap T^{*}_{b}\Omega\subset\mathscr{R}_{0}(\omega_{T})$ and the $x$-projection of $\mathscr{R}_{0}(\omega_{T})$ is included in $O^{T}$, $u_{1}=\psi(x,D_{x})u_{1}$ implies that $u_{1}$ is supported in $O^{T}$. Therefore, $u_{1}$ vanishes everywhere. Now, using (4.2), $\psi(x,D_{x})(u_{1,k})$ is bounded in $L^{2}(\Omega)$, so we also obtain by compactness that $(\psi(x,D_{x})u_{1,k})_{k\in{\mathbb{N}}}$ strongly converges to 00 in $H^{-1}(\Omega)$, thus getting a contradiction. ∎
@@ 5.1. Controllable $(\omega,\overline{\mathscr{O}})$ pairs
(5.1) $\left\{\begin{array}[]{ll}\partial_{t}^{2}y-\Delta_{A}y=v1_{\omega}&\text{in }(0,T)\times\Omega,\\ y(t,\cdot)=0&\text{on }(0,T)\times\partial\Omega,\\ (y(0,\cdot),\partial_{t}y(0,\cdot))=(0,0)&\text{in }\Omega.\end{array}\right.$
@@
(5.3) $||v||_{L^{2}(\omega_{T})}\leqslant C\|(y_{0}^{T},y_{1}^{T})\|_{H^{1}_{0}(\Omega)\times L^{2}(\Omega)}.$
@@
(5.5) $\left\{\begin{array}[]{ll}\partial_{t}^{2}u-\Delta_{A}u=0&\text{in }\,\,\mathcal{L}\\ u(t,.)=0&\text{on }\,\,\partial\mathcal{L}\\ (u(T,\cdot),\partial_{t}u(T,\cdot))=(u_{0}^{T},u_{1}^{T}),&\end{array}\right.$
@@
(5.9) $||v||_{L^{2}(\omega_{T})}+\|(y(T),\partial_{t}y(T))-\chi(y_{0}^{T},y_{1}^{T})\|_{H^{2}\cap H^{1}_{0}(\Omega)\times H^{1}_{0}(\Omega)}\leqslant C\|(y_{0}^{T},y_{1}^{T})\|_{H^{1}_{0}(\Omega)\times L^{2}(\Omega)}.$
@@
[…]
@@ 5.2. Pseudodifferential control when unique continuation holds
(5.12) $\widetilde{\mathscr{R}_{0}(\omega_{T})}=\Big{\{}(x,\xi)\in T_{b}^{*}\Omega\backslash 0,\text{ such that bicharacteristics $\gamma_{\rho}$}\\ \text{ issued from $(x,\xi)$ at time $T$ satisfy }\gamma_{\rho}({\mathbb{R}})\cap T^{*}(\omega_{T})\neq\emptyset\Big{\}}.$
@@
where $-\Delta_{A}$ is the operator $-{\text{div}}(A\nabla\cdot)$ in $\Omega$ with domain $H^{1}_{0}(\Omega)$ on $H^{-1}(\Omega)$.
@@
(5.15) $-{\text{div}}(A\nabla(y(T,\cdot)-\psi(x,D_{x})^{\star}y_{0}^{T}))=-(I-\psi(x,D_{x}))^{\star}(I-\psi(x,D_{x}))U_{0}^{T},$ in $\text{in }\Omega,$
(5.16) $\partial_{t}y(T,\cdot)-\psi(x,D_{x})^{\star}y_{1}=-(I-\psi(x,D_{x}))^{\star}(-\Delta_{A})^{-1}(I-\psi(x,D_{x}))U_{1}^{T},$ in $\text{in }\Omega.$
@@
One may wonder why this approach was not used in the proof of Theorem 5.1. The reason lies in the structure of formulas (5.15)–(5.16), which involve commutators with the operator $-\Delta_{A}$. While these commutators do not affect the regularity of the solutions, they significantly alter their support properties—particularly in the case of formula (5.15). As a result, this method is not well-suited for establishing Theorem 5.1.
@@ 6.2. Other observability techniques
Other than the microlocal tools employed in this article, the observability of waves has been often addressed employing multiplier methods [12] or Carleman estimates (see for instance [5]). Although they allow to refine global observability estimates when imposing conditions on the support of the initial data (see [12, Chapitre I, Section 9]) by reducing the observability time, these methods do not allow to get the sharp microlocal results in this paper.

### v2

@@ 3.8. A $1$-d example
[…]
@@
[…]
In this case, the bicharacteristic rays are particularly simple: they are the curves $t\mapsto x_{0}+t$ and $t\mapsto x_{0}-t$ for $x_{0}\in(-10,10)$ while these curves stay in the domain, bouncing back when meeting the boundary. Accordingly, in $1$-d, we can identify the characteristic manifold $\text{Char}(P_{A})$ with ${\mathbb{R}}\times\Omega\times\{-1,1\}$, depending if $\tau=|\xi|$, corresponding to $\epsilon=1$, or $\tau=-|\xi|$ corresponding to $\epsilon=-1$.
@@
$\mathscr{R}(\omega_{T})^{+}$ $=\{(t,x,+1)\text{ s.t. }t\in[0,T]\text{ and }-5+t<x<2+t\}$
$\mathscr{R}(\omega_{T})^{-}$ $=\{(t,x,-1)\text{ s.t. }t\in[0,T]\text{ and }-2-t<x<5-t\},$
@@ 4. Further observability results
The aim of this section is to refine Theorem 1.1 by analysing which microlocal components of the initial data of general wave solutions can be effectively observed from measurements taken on $\omega_{T}$.
@@ 4.1. Statement of the results
(4.7) $\|\psi(x)(u_{0},u_{1})\|_{H^{1}_{0}(\Omega)\times L^{2}(\Omega)}\leqslant C\|\partial_{t}u\|_{L^{2}(\omega_{T})}+C\|(1-\psi(x))(u_{0},u_{1})\|_{L^{2}(\Omega)\times H^{-1}(\Omega)}.$
Similarly, for any initial data $(u_{0},u_{1})\in L^{2}(\Omega)\times H^{-1}(\Omega)$,
(4.8) $\|\psi(x)(u_{0},u_{1})\|_{L^{2}(\Omega)\times H^{-1}(\Omega)}\leqslant C\|u\|_{L^{2}(\omega_{T})}+C\|(1-\psi(x))(u_{0},u_{1})\|_{H^{-1}(\Omega)\times(H^{2}\cap H^{1}_{0})^{\prime}(\Omega)}.$
@@
  • The constants appearing in Theorem 4.1 depend on both the observation time $T$ and the underlying metric $(a^{ij}(x))$. The dependence on the metric is very subtle and difficult to quantify explicitly, as it involves a large number of derivatives of the coefficients $a^{ij}$. This dependence can, in principle, be traced through a careful analysis of the proof of Theorem 3.5 in [16]. However, it is not readily expressible in closed form, since the operator norm of a pseudo-differential operator typically involves multiple derivatives of its symbol, making the exact dependence implicit and technically intricate.
@@
To illustrate the difficulty, consider the inequality (4.1). Let $\psi=\psi(t,x,D_{t},D_{x})\in\mathcal{A}^{0}$ be a pseudodifferential operator with $\operatorname{supp}\psi\cap T^{*}_{b}\mathcal{L}\subset\mathscr{R}(\omega_{T})$, and define the set
@@
By (4.2), any $u\in\mathcal{N}_{\psi}$ satisfies $u=\psi u\in H^{1}_{\text{loc}}(\mathcal{L})$, implying that $\mathcal{N}_{\psi}$ is compact and hence finite-dimensional.
Now consider $v=\partial_{t}u$ for $u\in\mathcal{N}_{\psi}$. Clearly, $v$ also solves the wave equation and satisfies $\partial_{t}v=0$ in $\omega_{T}$. However, there is no guarantee that v belongs to $\mathcal{N}_{\psi}$, as we do not have $(I-\psi)v=0$. In fact, since $\psi v=\psi(\partial_{t}u)$ and $v=\partial_{t}u$, their difference involves the commutator $\psi,\partial_{t}]u$, which does not vanish in general. Therefore, the naive use of the operator $\partial_{t}$ does not yield an operator acting invariantly on $\mathcal{N}_{\psi}$.
We have not been able to further analyze the structure of the sets $\mathcal{N}_{\psi}$. Whether or not $\mathcal{N}_{\psi}$ is non-trivial remains an open problem.
@@
Another example corresponds to the case in which $\Omega$ is the unit square, $A=Id$, and $\omega$ is an $\varepsilon$(>0>0)-neighborhood of the whole boundary. Note that, even if the square is not a smooth bounded domain and its boundary has tangencies of infinite order, this is not an impediment for our results to apply since the observation is made on a neighborhood of the whole boundary and, therefore, boundary phenomena are irrelevant. In this case, the GCC holds as soon as $T>\sqrt{2} (1-2\varepsilon)$ while unique continuation holds as soon as $T>(1-2\varepsilon)$. Therefore, when $T$ belongs to the intermediate interval $((1-2\varepsilon),\sqrt{2}(1-2\varepsilon))$, Theorem 4.7 applies, but the classical observability inequality based on GCC does not hold.
@@
Then all the items (1a), (2a), (1b) and (2b) in Theorem 4.7 hold for solutions $u$ of the wave equation (1.1) corresponding to initial data $(u_{0},u_{1})$ with either $u_{0}=0$ or $u_{1}=0$.
@@
Figure 4. When $\Omega={\mathbb{R}}^{d}$, $\omega$ is the unit ball and $\Delta_{A}$ is the flat Laplacian, the set $\mathcal{O}(\omega_{T})$, when projected in the physical space is supported in the ball of size $T$. In fact, in this case, as the bicharacteristics in this setting correspond to straight lines, O(ω_T) = {(x, ξ) ∈T_b^*Ω\0 s.t. ∃t ∈[0,T] satisfying x + ξt ∈ω}.
@@ 4.2. Proofs
By contradiction and following the proof of the estimate (4.11) of Theorem 4.7, we get a sequence $u_{1,k}$ such that $((I-\psi(x,D_{x}))u_{1,k})_{k\in{\mathbb{N}}}$ goes to 0 in $H^{-1}(\Omega)$, $\psi(x,D_{x})u_{1,k}$ is of unit norm in $H^{-1}(\Omega)$, and such that the corresponding solutions $u_{k}$ of (1.1) satisfies that $(\partial_{t}u_{k})_{k\in{\mathbb{N}}}$ goes to 0 in $L^{2}(\omega_{T})$. Consequently, up to a subsequence, we get $u_{1}$ such that $(u_{1,k})$ converges weakly to $u_{1}$ in $H^{-1}(\Omega)$, and $(I-\psi(x,D_{x}))u_{1}=0$, and such that the corresponding solution of (1.1) with initial data $(0,u_{1})$ satisfies $\partial_{t}u=0$ in $(0,T)\times\omega$. By the above uniqueness result, we thus get that $u_{1}=0$ in the set $O^{T}=\{x\in\overline{\Omega},\,d(x,\omega)<T\}$. Finally, since $\psi=\psi(x,D_{x})\in\mathcal{B}^{0}$ satisfies $\operatorname{Supp}(\psi)\cap T^{*}_{b}\Omega\subset\mathscr{R}_{0}(\omega_{T})$ and the $x$-projection of $\mathscr{R}_{0}(\omega_{T})$ is included in $O^{T}$, $u_{1}=\psi(x,D_{x})u_{1}$ implies that $u_{1}$ is supported in $O^{T}$. Therefore, $u_{1}$ vanishes everywhere. Now, using (4.2), $\psi(x,D_{x})(u_{1,k})$ is bounded in $L^{2}(\Omega)$, so we also obtain by compactness that $(\psi(x,D_{x})u_{1,k})_{k\in{\mathbb{N}}}$ strongly converges to 0 in $H^{-1}(\Omega)$, thus getting a contradiction. ∎
@@ 5.1. Controllable $(\omega,\overline{\mathscr{O}})$ pairs
(5.1) $\left\{\begin{array}[]{ll}\partial_{t}^{2}y-\Delta_{A}y=v1_{\omega}\quad&\text{in }(0,T)\times\Omega,\\ y(t,\cdot)=0\quad&\text{on }(0,T)\times\partial\Omega,\\ (y(0,\cdot),\partial_{t}y(0,\cdot))=(0,0)&\text{in }\Omega.\end{array}\right$.
@@
(5.3) $v||_{L^{2}(\omega_{T})}\leqslant C\|(y_{0}^{T},y_{1}^{T})\|_{H^{1}_{0}(\Omega)\times L^{2}(\Omega)}.$
@@
[…]
@@
(5.9) $v||_{L^{2}(\omega_{T})}+\|(y(T),\partial_{t}y(T))-\chi(y_{0}^{T},y_{1}^{T})\|_{H^{2}\cap H^{1}_{0}(\Omega)\times H^{1}_{0}(\Omega)}\leqslant C\|(y_{0}^{T},y_{1}^{T})\|_{H^{1}_{0}(\Omega)\times L^{2}(\Omega)}.$
@@
The observability estimate (4.8) easily provides the coercivity and strict convexity of the functional $J_{\chi}$ on the space $X_{obs}=\overline{L^{2}(\Omega)\times H^{-1}(\Omega)}^{\|\cdot\|_{obs}}$, where the norm $\| \cdot\|_{obs}$ is given by
@@ 5.2. Pseudodifferential control when unique continuation holds
[…]
@@
where $\Delta_{A}$ is the operator ${\text{div}}(A\nabla\cdot)$ in $\Omega$ with domain $H^{1}_{0}(\Omega)$ on $H^{-1}(\Omega)$.
@@
(5.15) $-{\text{div}}(A\nabla(y(T,\cdot)-\psi(x,D_{x})^{\star}y_{0}^{T}))=-(I-\psi(x,D_{x}))^{\star}(I-\psi(x,D_{x}))U_{0}^{T},$ $\text{in }\Omega,$
(5.16) $\partial_{t}y(T,\cdot)-\psi(x,D_{x})^{\star}y_{1}=-(I-\psi(x,D_{x}))^{\star}(-\Delta_{A})^{-1}(I-\psi(x,D_{x}))U_{1}^{T},$ $\text{in }\Omega.$
@@
One may wonder why this approach was not used in the proof of Theorem 5.1. The reason lies in the structure of formulas (5.15)–(5.16), which involve commutators with the operator $\Delta_{A}$. While these commutators do not affect the regularity of the solutions, they significantly alter their support properties—particularly in the case of formula (5.15). As a result, this method is not well-suited for establishing Theorem 5.1.
@@ 6.2. Other observability techniques
Other than the microlocal tools employed in this article, the observability of waves has been often addressed employing multiplier methods [13] or Carleman estimates (see for instance [6]). Although they allow to refine global observability estimates when imposing conditions on the support of the initial data (see [13, Chapitre I, Section 9]) by reducing the observability time, these methods do not allow to get the sharp microlocal results in this paper.